OpenStudy (anonymous):

If (2 − 3i) + (x + yi) = 6, what is x + yi?

OpenStudy (zarkon):

\[ (2 − 3i) + (x + yi) = 6\] \[(2+x)+(-3+y)i=6+0i\]

OpenStudy (mathmale):

Beginning with the given \[(2 − 3i) + (x + yi) = 6\] eliminate the parentheses. Rearrange the terms so that you have only x+yi on the left side.